OpenStudy (crashonce):

Prove that if 2cosBsinC = sinA then ABC is an isosceles triangle

OpenStudy (crashonce):

@iGreen @ganeshie8

OpenStudy (asnaseer):

if ABC is indeed an isosceles triangle then \(B=C\) and \(A=\pi-(B+C)=\pi-2B\), therefore:\[2\cos(B)\sin(C)=2\cos(B)\sin(B)=\sin(2B)\]Hopefully you can complete the proof from here...